A fisherman takes his boat out to sea and then back to the harbor. The entire trip lasts 8 hours. If he travels out at 18 mph and follows the same route back at 14 mph, how many miles does he travel altogether?
Answer by pushpaharan(47)

lets assume distance travelled = x
as you know speed = distance /time
total time taken = x/18 + x/14 = 8
7x/126 + 9x/126 = 8
16x/126 = 8
x = 8*126/16 = 63 miles